Note:

When I use the word "bar" after the complex number expressions, I simply mean there is a bar over each expression.

Given z = 3 - 4i and w = 8 + 3i, write each expression in the standard form
a + bi.

1. z + z bar

2. w - w bar

Answer by ikleyn(52775)

.
Given z = 3 - 4i and w = 8 + 3i, write each expression in the standard form
a + bi.
1. z + z bar
2. w - w bar
~~~~~~~~~~~~~~~~~

(1)  z + z bar = (3-4i) + (3+4i) = 3 + 3 = 6.

     Post-solution note :  z + z bar is always doubled real part of z,
                           and has zero imaginary part.



(2)  w - w bar = (8+3i) - (8-3i) = 3i - (-3i) = 3i + 3i = 6i.

     Post-solution note :  w - w bar is always doubled imaginary part of w  
                           and has zero real part.

Solved.